You own a workshop that produces custom handmade frames for diplomas. You are negotiating a contract with Loyola to sell the frames at the bookstore. Loyola decided to offer a trial contract where they will stock 20 frames at the bookstore. You estimate that the material costs for each frame is $15. You also estimate that it takes your employee 15 labor hours to produce the first frame. You pay your employee $25 per hour. Based on your observations, you noticed that your employee is on an 85% learning curve.

What is the lowest total bid you should submit if you charge a 9.6% markup, if you account for learning curve?

What is the lowest total bid you should submit if you charge a 9.6% markup, if you account for learning curve?
Answer #1

As per Wright's model, Y = aXb

Y = the cumulative average time per unit.
X = the cumulative number of units produce = 20
a = time (or cost) required to produce the first unit = 15
b = slope of the function when plotted on log-log paper.
   = log of the learning rate/log of 2.

Y = 15 * 20 ^ (log 85%)/(log2)

Y = 15*20^- 0.234 = 7.441

This is average time, hence time for 20 frames = 20 * 7.441 = 148.82 hours

Labor cost = 25 * 148.82 = $3720.5

Material cost = 20*15 = $300

Total cost = $3720.5 + 300 = $4020.5

Bid price = Total cost + Markup% * Total cost = $4406.5

(Note - Markup is percentage of cost and Margin is %age of sale price)